# Improved Extended Regular Expression Matching

### v1

@@ 4.1 Clustering
Let $R$ be an extended regular expression of length $m$ containing $k$ extended operators. We will use a clustering of the parse tree $T(R)$ into $O(k)$ node-disjoint clusters such that each cluster contains at most one node labeled with an extended operator. Similar clustering techniques (also called modules) are used in previous solutions [56, 57].
We define the set of extended nodes $P$ to be the subset of nodes $v$ in $T(R)$ such that (i) $v$ is labeled by an extended operator or (ii) $v$ is the lowest common ancestor of two nodes that are labeled by an extended operator. The edges in $T(R)$ from a node $p$ in $P$ to any children of $p$ are external edges and all other edges are internal edges. Deleting all external edges partitions $T(R)$ into node-disjoint subtrees, that we call clusters, and the set of all these clusters is the cluster partition $CS$ of $T(R)$. Contracting all internal edges induces a macro tree, where each cluster is represented by a single node. We will use standard tree terminology on the macro tree and refer to clusters as child clusters, internal clusters, parent clusters, child clusters, etc. We can construct the clustering in a bottom-up traversal of $T(R)$ in $O(m)$ time. See Figure 4 for an illustration.
Figure 4: In $a)$ is shown the clustering of the parse tree of an extended regular expression. The dark gray nodes are the nodes from $P$. In $b)$ is the automaton corresponding to cluster $C_{3}$ in Figure $a)$.
@@
Let $P$ be the set of extended nodes in $T(R)$ and let $CS$ be the corresponding cluster partition. (i) By definition a leaf cluster in $CS$ cannot contain an extended node $p\in P$. (ii) Consider an internal cluster $C\in CS$. By definition, an internal cluster must contain at least one extended node. Note that an extended node cannot be in the same cluster as any of its descendants. Assume for contradiction that $C$ contains two distinct extended nodes $p_{1},p_{2}\in P$. Then, the lowest common ancestor $p_{3}=\mathsf{lca}(p_{1},p_{2})$ is an extended node also in $C$ (if $p_{3}$ was not in $C$, then both $p_{1}$ and $p_{2}$ can’t be in $C$). If $p_{3}\neq p_{1}$, $p_{3}$ is a proper ancestor of $p_{1}$ and thus $p_{3}$ cannot be in $C$. Symmetrically, if $p_{3}\neq p_{2}$, $p_{3}$ can also not be in $C$. (iii) Consider the macro tree $M$. Then, $M$ has at most $2k$ leaf clusters, since the parents of the roots of the leaf clusters are all labeled by an extended operator. It follows that there are at most 2k12k-1 internal clusters. In total, we have at most 4k14k-1 clusters. ∎
@@ 4.2 General Match Graphs
We will need a generalization of our match graphs from Section 3.1 along with efficient operations on them to implement our algorithm. Given a TNFA $A$ and a pair of states $(s,t)$, the match graph for $A$, $s$, and $t$, denoted $G(A,s,t)$, is the directed graph on $n+1$ vertices $g_{1},\ldots g_{n+1}$ such that there is an edge from $g_{i}$ to $g_{j}$ iff there is a path from $s$ to $t$ in $A$ matching Q[i,j1]Q[i,j-1], $1\leq i\leq j\leq n+1$. As before, if there is a path from $s$ to $t$ matching the empty string, every vertex in $G(A,s,t)$ has a self-loop. If $s=t$, then every vertex in $G(A,s,t)$ also has a self-loop.
@@
  • For $G^{*}$, we add an edge $(g_{i},g_{j})$ for each pair $i,j$ with a directed path from $g_{i}$ to $g_{j}$ in $G(v)$ and add self-loops to all vertices.
@@ 4.3 Algorithm
We construct the parse tree $T(R)$ and the clustering according to Lemma 2. To speed up the dynamic programming algorithm of Section 3, we will show how to efficiently compute the match graph of the root of each cluster in a bottom-up traversal of the macro tree. At a cluster $C\in CS$ with root $v$, construct the match graph $G(v)$ according to the following cases.
Case 1: $C$ is a leaf cluster.
We construct the TNFA $A_{C}$ and compute the match graph $G(v)=G(A_{C},\theta,\phi)$ according to Lemma 4.
Case 2: $C$ is an internal cluster.
Let $p\in P$ be the single extended node in $C$. We compute $G(v)$ as follows.
Step 1:
Compute $G(p)$. There are two cases.
@@
From the clustering, it follows that $u$ is labeled $\neg$. We compute $G(p)$ from $G(u)$ as described in Lemma  3.
@@
Figure 5: Illustration of the matching paths represented by the four match graphs computed by simulating $A_{C}$.
@@ 4.4 Correctness
We will show that Q[i,j1]Q[i,j-1] matches $R(v)$ if and only if $(g_{i},g_{j})\in G$. We first define an automaton $\widehat{A}_{C}$ used only for the proof. The automaton $\widehat{A}_{C}$ contains the same states and transitions as the TNFA $A_{C}$, in addition to the following paths from $\theta^{\prime}$ to $\phi^{\prime}$: For each interval [i,j1][i,j-1] of $Q$ that matches $R(p)$, we add a path from $\theta^{\prime}$ to $\phi^{\prime}$ with transition labels concatenating to Q[i,j1]Q[i,j-1]. Thus, $\widehat{A}_{C}$ accepts a string if and only if it matches $R(v)$. Note, that $A_{C}$ is a subgraph of $\widehat{A}_{C}$.
First, we show that Q[i,j1]Q[i,j-1] matches $R(v)$ implies $(g_{i},g_{j})\in G$. Let $S$ be the matching path of Q[i,j1]Q[i,j-1] from $\theta$ to $\phi$ in $\widehat{A}_{C}$.
Case 1: If $\theta^{\prime}\not\in S$, then $S$ is contained in $A_{C}$, and thus $(g_{i},g_{j})\in G_{\theta,\phi}$ by the construction of $G_{\theta,\phi}$.
Case 2: Otherwise, we decompose $S$, splitting it into subpaths $S_{0},...,S_{l}$ at every occurrence of $\theta^{\prime}$ and $\phi^{\prime}$. Note that if $\theta=\theta^{\prime}$, then subpath $S_{0}$ is empty, and likewise, if $\phi=\phi^{\prime}$, then $S_{l}$ is empty. This decomposition of the path $S$ gives rise to a corresponding decomposition of the string $Q$ into $l+1$ substrings:
$Q[i,j-1]=Q[i_{0},i_{1}-1]Q[i_{1},i_{2}-1]...Q[i_{l},i_{l+1}-1],$
where $i_{0}=i$ and $i_{l+1}=j$. Note, that $Q[i_{h},i_{h+1}-1]$ is the empty string, if either the subpath $S_{h}$ is empty (can happen for $h\in\{0,l\}$) or $S$ only contains transitions labeled with $\epsilon$.
Since $S_{0}$ is a path in $A_{C}$ from $\theta$ to $\theta^{\prime}$ then $(g_{i_{0}},g_{i_{1}})\in G_{\theta,\theta^{\prime}}$. Similarly, since $S_{l}$ is a path in $A_{C}$ from $\phi^{\prime}$ to $\phi$ then $(g_{i_{l}},g_{i_{l+1}})\in G_{\phi^{\prime},\phi}$. For $0<h<l$, if $h$ is even, then $S_{h}$ is a path from $\phi^{\prime}$ to $\theta^{\prime}$, in which case, $(g_{i_{h}},g_{i_{h+1}})\in G_{\phi^{\prime},\theta^{\prime}}$. Similarly, if $h$ is odd, then $S_{h}$ is a path from $\theta^{\prime}$ to $\phi^{\prime}$, in which case $(g_{i_{h}},g_{i_{h+1}})\in G(p)$ by the correctness of $G(p)$. Therefore, for even $h$, $0<h<l$, there is an edge $(g_{i_{h}},g_{i_{h+1}})\in G_{\phi^{\prime},\theta^{\prime}}$ and an edge $(g_{i_{h+1}},g_{i_{h+2}})\in G(p)$. This implies $(g_{i_{h}},g_{i_{h+2}})\in G_{\phi^{\prime},\theta^{\prime}}\odot G(p)$, which in turn implies that $(g_{i_{2}},g_{i_{l}})\in(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}$. It follows that, $(g_{i},g_{j})=(g_{i_{0}},g_{i_{l+1}})\in G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$.
@@
Combining case 1 and 2, we get that $(g_{i},g_{j})\in G_{\theta,\phi}\mid(G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi})=G$. We now show that $(g_{i},g_{j})\in G$ implies that Q[i,j1]Q[i,j-1] matches $R(v)$. If $(g_{i},g_{j})\in G$, then the edge $(i,j)$ is either in $G_{\theta,\phi}$, in $G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$ or in both.
Case 1: If $(g_{i},g_{j})\in G_{\theta,\phi}$, then there is a path in $A_{C}$ from $\theta$ to $\phi$ matching Q[i,j1]Q[i,j-1], so Q[i,j1]Q[i,j-1] matches $R(v)$.
@@
Case 2: Otherwise, $(g_{i},g_{j})\in G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$. Then there exist indices $i_{0},...,i_{l+1}$ with $i_{0}=i$ and $i_{l+1}=j$ such that $(g_{i_{0}},g_{i_{1}})$ is an edge in $G_{\theta,\theta^{\prime}}$, $(g_{i_{l}},g_{i_{l+1}})$ is an edge in $G_{\phi^{\prime},\phi}$, $(g_{i_{1}},g_{i_{2}})$ is an edge in $G(p)$ and $(g_{i_{2}},g_{i_{l}})$ is an edge in $(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}$. This implies that for even $h$, with $0<h<l$, the edge $(g_{i_{h}},g_{i_{h+2}})$ is in $G_{\phi^{\prime},\theta^{\prime}}\odot G(p)$. Then $(g_{i_{h}},g_{i_{h+1}})\in G_{\phi^{\prime},\theta^{\prime}}$ and $(g_{i_{h+1}},g_{i_{h+2}})\in G(p)$. Each of the edges $(g_{i_{h}},g_{i_{h+1}})$ for $0\leq h\leq l+1$ corresponds to a path in $\widehat{A}_{C}$ matching $Q[i_{h},i_{h+i}-1]$, and the concatenation of these paths is a path from $\theta$ to $\phi$ in $\widehat{A}_{C}$, that matches Q[i,j1]Q[i,j-1]. ∎

### v2

@@ 4.1 Clustering
Let $R$ be an extended regular expression of length $m$ containing $k$ extended operators. We will use a clustering of the parse tree $T(R)$ into $O(k)$ node-disjoint clusters such that each cluster contains at most one node labeled with an extended operator. While clustering techniques (also called modules) are also in previous solutions [66, 67], our use of clustering in combination with match graphs and finite automaton simulation is new and differs significantly from previous approaches.
We define the set of extended nodes $P$ to be the subset of nodes $v$ in $T(R)$ such that (i) $v$ is labeled by an extended operator or (ii) $v$ is the lowest common ancestor of two nodes that are labeled by an extended operator. The edges in $T(R)$ from a node $p$ in $P$ to any children of $p$ are external edges and all other edges are internal edges. Deleting all external edges partitions $T(R)$ into node-disjoint subtrees that we call clusters, and the set of all these clusters is the cluster partition $CS$ of $T(R)$. Contracting all internal edges induces a macro tree, where each cluster is represented by a single node. We will use standard tree terminology on the macro tree and refer to clusters as child clusters, internal clusters, parent clusters, child clusters, etc. We can construct the clustering in a bottom-up traversal of $T(R)$ in $O(m)$ time. See Figure 4 for an illustration.
Figure 4: $(a)$ The clustering of the parse tree of an extended regular expression. The dark gray nodes are the nodes from $P$. $(b)$ The automaton corresponding to cluster $C_{3}$ in $(a)$.
@@
Let $P$ be the set of extended nodes in $T(R)$ and let $CS$ be the corresponding cluster partition. (i) By definition, a leaf cluster in $CS$ cannot contain an extended node $p\in P$. (ii) Consider an internal cluster $C\in CS$. By definition, an internal cluster must contain at least one extended node. Note that an extended node cannot be in the same cluster as any of its descendants. Assume for contradiction that $C$ contains two distinct extended nodes $p_{1},p_{2}\in P$. Then, the lowest common ancestor $p_{3}=\mathsf{lca}(p_{1},p_{2})$ is an extended node also in $C$ (if $p_{3}$ was not in $C$, then both $p_{1}$ and $p_{2}$ can’t be in $C$). If $p_{3}\neq p_{1}$, $p_{3}$ is a proper ancestor of $p_{1}$ and thus $p_{3}$ cannot be in $C$. Symmetrically, if $p_{3}\neq p_{2}$, $p_{3}$ can also not be in $C$. (iii) Consider the macro tree $M$. Then, $M$ has at most $2k$ leaf clusters, since the parents of the roots of the leaf clusters are all labeled by an extended operator. It follows that there are at most 2k12k-1 internal clusters. In total, we have at most 4k14k-1 clusters. ∎
@@
Intuitively, the extended transition represents the extended regular expression $R(p)$. Let $Q$ be a string that matches the regular expression $R(v)$. Then, there is a decomposition of $Q$ into substrings $Q=Q_{1}\odot\tilde{Q}_{1}\odot Q_{2}\odot\tilde{Q}_{2}\odot\cdots\odot\tilde{Q}_{\ell-1}\odot Q_{\ell}$, such that $\tilde{Q}_{i}$ matches the extended regular expression $R(p)$ and $Q_{i}$ matches parts of the non-extended part of the regular expression $R(v)$ for all $i$. If we replace each $\tilde{Q}_{i}$ with $\beta$, we obtain a string $Q^{\prime}=Q_{1}\odot\beta\odot Q_{2}\odot\beta\odot\cdots\odot\beta\odot Q_{\ell}$ that is accepted by $A_{C}$.
Example 1.
Let $v$ be the root of cluster $C_{3}$ in Figure 4(a) and let $Q=cabcbba$. There are several ways to match $Q$ to the regular expression $R(v)$. One such way is to match the two substrings $Q[2,3]=ab$ and $Q[5,7]=bb$ to $R(p)=\neg(a\mid b)$. Replacing these two substrings with $\beta$ we get $Q^{\prime}=c\beta c\beta a$ which is accepted by the automaton in Figure 4(b).
@@ 4.2 General Match Graphs
In this section we generalize our match graphs from Section 3.1 and give efficient operations on them to implement our algorithm. Given a TNFA $A$ and a pair of states $(s,t)$, the match graph for $A$, $s$, and $t$, denoted $G(A,s,t)$, is the directed graph on $n+1$ vertices $g_{1},\ldots g_{n+1}$ such that there is an edge from $g_{i}$ to $g_{j}$ iff there is a path from $s$ to $t$ in $A$ matching Q[i,j1]Q[i,j-1], $1\leq i\leq j\leq n+1$. As before, if there is a path from $s$ to $t$ matching the empty string, every vertex in $G(A,s,t)$ has a self-loop. If $s=t$, then every vertex in $G(A,s,t)$ also has a self-loop.
@@
  • For $G^{*}$, we add an edge $(g_{i},g_{j})$ for each pair $i,j$ with a directed path from $g_{i}$ to $g_{j}$ in $G$ and add self-loops to all vertices.
@@ 4.3 Algorithm
We are now ready to describe the improved algorithm. We first construct the parse tree $T(R)$ and the clustering according to Lemma 2. To speed up the dynamic programming algorithm of Section 3, we will show how to efficiently compute the match graph of the root of each cluster in a bottom-up traversal of the macro tree.
The main idea is the following. As described in Section 4.1 if a string $Q$ matches $R(v)$ then $Q$ can be partitioned into substrings $Q=Q_{1}\odot\tilde{Q}_{1}\odot Q_{2}\odot\tilde{Q}_{2}\odot\cdots\odot\tilde{Q}_{\ell-1}\odot Q_{\ell}$, such that $\tilde{Q}_{i}$ matches $R(p)$ for all $i$, and the string $Q^{\prime}=Q_{1}\odot\beta\odot Q_{2}\odot\beta\odot\cdots\odot Q_{\ell-1}\odot\beta\odot Q_{\ell}$ is accepted by $A_{C}$. Furthermore, the string $Q_{1}$ must match a path from $\theta$ to $\theta^{\prime}$ in $A_{C}$, $Q_{\ell}$ must match a path from $\phi^{\prime}$ to $\phi$, and for all other $i$, $Q_{i}$ must match a path from $\phi^{\prime}$ to $\theta^{\prime}$.
Example 2.
Consider the example from Example 1 and the automaton from Figure 4(b). Here $Q^{\prime}=c\beta c\beta a$ and $Q_{1}=c$ matches a path from $\theta$ to $\theta^{\prime}$, $Q_{2}=c$ matches a path from $\phi^{\prime}$ to $\theta^{\prime}$, and $Q_{3}=a$ matches a path from $\theta^{\prime}$ to $\theta$.
We therefore compute the match graphs $G(A_{C},\theta,\theta^{\prime}),G(A_{C},\phi^{\prime},\theta^{\prime})$, and $G(A_{C},\phi^{\prime},\phi)$. To include the case where the parsing of $Q$ does not use the subexpression $R(p)$ we also compute the match graph $G(A_{C},\theta,\phi)$. We then combine these match graphs with the match graph $G(p)$ using the match graph operations from the previous section which allows us to write the combinations of the match graphs as a regular expression of match graphs (see Figure 5(a)). The intuition is that there will be an edge $(l,r)$ in the match graph $G(v)$ if there is $(l,r)$ in is an edge in $G(A_{C},\theta,\phi)$, or if there exists a sequence of edges $(l_{1},r_{1}),(\tilde{l}_{1},\tilde{r}_{1}),(l_{2},r_{2}),(\tilde{l}_{2},\tilde{r}_{2}),\ldots,(\tilde{l}_{\ell-1},\tilde{r}_{\ell-1}),(l_{\ell},r_{\ell})$ such that $l_{1}=l$, $r_{\ell}=r$, $\tilde{l}_{i}=r_{i}$ for all $i$, and $l_{i}=\tilde{r}_{i-1}$ for all $i\neq 1$. Furthermore, $(l_{1},j_{1})\in G(A_{C},\theta,\theta^{\prime})$, $(l_{\ell},r_{\ell})\in G(A_{C},\phi^{\prime},\phi)$, $(\tilde{l}_{i},\tilde{r}_{i})\in G(p)$ and $(l_{i},r_{i})\in G(A_{C},\phi^{\prime},\theta^{\prime})$ for $i\not\in\{1,\ell\}$. See Figure 5(b) for an example.
We now present the full algorithm. First, construct the parse tree $T(R)$, the clustering according to Lemma 2, the corresponding macro tree $M$ of the clusters $CS$, and the set of extended nodes $P$. We traverse the macro tree bottom-up. At each cluster $C$ with root $v$, we compute the match graph $G(v)$ as follows. If $C$ is a leaf cluster, $C$ contains no extended node. We directly construct the TNFA $A_{C}$ and compute the match graph $G(v)=G(A_{C},\theta,\phi)$ according to Lemma 4. Otherwise, $C$ is an internal cluster with a single extended node $p\in P$. We perform the following steps.
Step 1: Compute $G(p)$.
We first compute $G(p)$. Since we compute the match graphs of the clusters bottom up we have already computed the match graphs of $p$’s children in the parse tree, as these will be the roots of the child clusters of $C$ in the macro tree. There are two cases.
@@
From the clustering, it follows that $u$ is labeled $\neg$. We compute $G(p)$ from $G(u)$ as described in Lemma 3.
@@
Figure 5: $(a)$ The regular expression of the four match graphs used in step 3 of the algorithm. $(b)$ The match graphs computed for cluster $C_{3}$ in Figure 4 and the string $Q=cabcbba$ from Example 1 and 2.
@@ 4.4 Correctness
We will show that $(g_{l},g_{r})\in G$ if and only if Q[l,r1]Q[l,r-1] matches $R(v)$. We first define an automaton $\widehat{A}_{C}$ used only for the proof. The automaton $\widehat{A}_{C}$ contains the same states and transitions as the TNFA $A_{C}$, in addition to the following paths from $\theta^{\prime}$ to $\phi^{\prime}$: For each interval [i,j1][i,j-1] of $Q$ that matches $R(p)$, we add a path from $\theta^{\prime}$ to $\phi^{\prime}$ with transition labels concatenating to Q[i,j1]Q[i,j-1]. Thus, $\widehat{A}_{C}$ accepts a string if and only if it matches $R(v)$. Note, that $A_{C}$ is a subgraph of $\widehat{A}_{C}$.
@@
First we show that $(g_{l},g_{r})\in G$ implies that Q[l,r1]Q[l,r-1] matches $R(v)$. Assume $(g_{l},g_{r})\in G$. Then $(g_{l},g_{r})$ is either in $G_{\theta,\phi}$, in $G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$ or in both. Suppose first that $(g_{l},g_{r})\in G_{\theta,\phi}$. Then there is a path in $A_{C}$ from $\theta$ to $\phi$ matching Q[l,r1]Q[l,r-1], so Q[l,r1]Q[l,r-1] matches $R(v)$. Otherwise, we have that $(g_{l},g_{r})\in G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$. Then there is a sequence of edges
$(g_{l_{1}},g_{r_{1}}),(g_{\tilde{l}_{1}},g_{\tilde{r}_{1}}),(g_{l_{2}},g_{r_{2}}),(g_{\tilde{l}_{2}},g_{\tilde{r}_{2}}),\ldots,(g_{\tilde{l}_{\ell-1}},g_{\tilde{r}_{\ell-1}}),(g_{l_{\ell}}g_{r_{\ell}})$
@@
where $l_{1}=l$, $r_{\ell}=r$, $\tilde{l}_{i}=r_{i}$ for all $i$, and $l_{i}=\tilde{r}_{i-1}$ for all $i\neq 1$, such that $(g_{l_{1}},g_{j_{1}})\in G(A_{C},\theta,\theta^{\prime})$, $(g_{l_{\ell}},g_{r_{\ell}})\in G(A_{C},\phi^{\prime},\phi)$, $(g_{\tilde{l}_{i}},g_{\tilde{r}_{i}})\in G(p)$ and $(g_{l_{i}},g_{r_{i}})\in G(A_{C},\phi^{\prime},\theta^{\prime})$ for $i\not\in\{1,\ell\}$. The edge $(g_{l_{1}},g_{r_{1}})$ corresponds to a path in $A_{C}$ matching $Q[l_{1},r_{1}-1]=Q[l,r_{1}-1]$ and the edge $(g_{l_{\ell}},g_{r_{\ell}})$ corresponds to a path in $A_{C}$ matching $Q[l_{\ell},r_{\ell}-1]=Q[l_{\ell},r-1]$. Each of the edges $(g_{l_{i}},g_{r_{i}})$, for $1<i<\ell$, corresponds to a path in $A_{C}$ matching $Q[l_{i},r_{i}-1]$, and each of the edges $(g_{\tilde{l}_{i}},g_{\tilde{r}_{i}})$, for $1\leq i<\ell$, corresponds to a path in $\widehat{A}_{C}$ from $\theta^{\prime}$ to $\phi^{\prime}$ matching $Q[\tilde{l}_{i},\tilde{r}_{i}-1]$. The concatenation of these paths is a path from $\theta$ to $\phi$ in $\widehat{A}_{C}$, that matches Q[l,r1]Q[l,r-1]. Thus Q[l,r1]Q[l,r-1] matches $R(v)$.
Next, we show that if Q[l,r1]Q[l,r-1] matches $R(v)$ then $(g_{l},g_{r})\in G$. Let $\Gamma$ be the matching path of Q[l,r1]Q[l,r-1] from $\theta$ to $\phi$ in $\widehat{A}_{C}$. Suppose that $\theta^{\prime}\not\in\Gamma$. Then $\Gamma$ is fully contained in $A_{C}$, and thus $(g_{l},g_{r})\in G_{\theta,\phi}$ by the construction of $G_{\theta,\phi}$. Otherwise, we decompose $\Gamma$, splitting it into subpaths
$\Gamma_{1},\tilde{\Gamma}_{1},\Gamma_{2},\tilde{\Gamma}_{2},\ldots,\tilde{\Gamma}_{\ell-1},\Gamma_{\ell}$
at every occurrence of $\theta^{\prime}$ and $\phi^{\prime}$. Note that if $\theta=\theta^{\prime}$, then subpath $\Gamma_{1}$ is empty, and likewise, if $\phi=\phi^{\prime}$, then $\Gamma_{\ell}$ is empty. This decomposition of the path $\Gamma$ corresponds to a decomposition of the string Q[l,r1]Q[l,r-1] into substrings:
$Q_{1},\tilde{Q}_{1},Q_{2},\tilde{Q}_{2},\ldots,\tilde{Q}_{\ell-1},Q_{\ell}$
such that $Q[l,r-1]=Q_{1}\odot\tilde{Q}_{1}\odot\cdots\odot\tilde{Q}_{\ell-1}\odot Q_{\ell}$. Note, that $Q_{i}$ (resp. $\tilde{Q}_{i}$) is the empty string, if either the subpath $\Gamma_{i}$ (resp. $\tilde{\Gamma}_{i}$) is empty (can happen for $\Gamma_{1}$ and $\Gamma_{\ell}$) or if $\Gamma$ only contains transitions labeled with $\epsilon$. Let $Q_{i}=Q[l_{i},r_{i}]$ and $\tilde{Q}_{i}=Q[\tilde{l}_{i},\tilde{r}_{i}]$. Since $\Gamma_{1}$ is a path in $A_{C}$ from $\theta$ to $\theta^{\prime}$ then $(g_{l_{1}},g_{r_{1}})\in G_{\theta,\theta^{\prime}}$. Similarly, since $\Gamma_{\ell}$ is a path in $A_{C}$ from $\phi^{\prime}$ to $\phi$ then $(g_{l_{\ell}},g_{r_{\ell}})\in G_{\phi^{\prime},\phi}$. For $1\leq i<\ell$, $\tilde{\Gamma}_{i}$ is a path from $\theta^{\prime}$ to $\phi^{\prime}$, in which case $(g_{\tilde{l}_{i}},g_{\tilde{r}_{i}})\in G(p)$. For $1<i<\ell$, $\Gamma_{i}$ is a path from $\phi^{\prime}$ to $\theta^{\prime}$, in which case, $(g_{l_{i}},g_{r_{i}})\in G_{\phi^{\prime},\theta^{\prime}}$. It follows that, $(g_{l},g_{r})=(g_{l_{1}},g_{r_{\ell}})\in G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi}$.
It follows that $(g_{l},g_{r})\in G_{\theta,\phi}\mid(G_{\theta,\theta^{\prime}}\odot G(p)\odot(G_{\phi^{\prime},\theta^{\prime}}\odot G(p))^{*}\odot G_{\phi^{\prime},\phi})=G$. ∎